OpenStudy (anonymous):

How do I simplify this? -3x/x^2-9 + 4/2x-6

OpenStudy (anonymous):

-3x/(x^2-9) + 4/(2x-6)

OpenStudy (anonymous):

\[\frac{ -3x }{ x ^{2} -9 } + \frac{ 4 }{ 2x-6 }\]

OpenStudy (anonymous):

-3x/(x^2-9) + 4/(2x-6) -3x/(x-3)(x+3) + 4/(2(x-3)) -3x/(x-3)(x+3) + 2/(x-3) multiply the second equation by (x+3)/(x+3) to get a common denominator -3x/(x-3)(x+3) + 2(x+3)/(x+3)(x-3) add the numerators. keep the common denominator [-3x + 2(x+3)]/[(x+3)(x-3)] [-3x + 2x + 6]/[(x+3)(x-3)] (-x+6)/[(x+3)(x-3)] \[\frac{-x+6}{(x^2-9)}\]

OpenStudy (anonymous):

Thank you very much! @tomo

OpenStudy (anonymous):

you're welcome